The dimension of a rectangle are such that the length is 5 inches more than its width. If the length were doubled and if the width were decreased by 2 inches, the are would be increased by 190inches squared. What are the length and width of the rectangle?
Answer by mananth(16946)

width = x
length = x+5
2(x+5) new length
(x-2) new width
2(x+5)(x-2)-x(x+5) =190
2(x^2+3x-10)-x^2-5x=190
2x^2+6x-20-x^2-5x=190
x^2+x-210=0
x^2+15x-14x-210=0
x(x+15)-14(x+15)=0
(x+15)(x-14)=0
x= 14 which is positive
width = 14 inches
length =19 inches